# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
Assume that $\delta$ is defined in (5). Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $|\tilde{\mathcal{C}}_{3}|\geq 2\sqrt{\delta}n$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
In the following of this section, we prove the above two theorems respectively.
5.1 Proof of Theorem 5.1
In this subsection, we give the proof of Theorem 5.1.
Proof of Theorem 5.1.
Suppose that $\mathcal{D}$ does not contain transversal directed Hamilton cycles. Let $\epsilon,\delta$ be defined in (5), and choose $\eta$ with $\delta\ll\eta\ll 1$. Denote $[m]:=(\tilde{\mathcal{C}}_{1}\cup\tilde{\mathcal{C}}_{2})\setminus I^{1}$ and $\mathcal{C}_{\rm bad}:=[m+1,n]$. By adding colors to $\mathcal{C}_{\rm bad}$ if necessary we may assume $m=(1-4\sqrt{\delta})n$. Recall that $\epsilon\ll\delta$. Hence for each $i\in[2,m]$, there exists a new partition $(A_{i},B_{i},L_{i})$ of $D_{i}$ such that
  • (C1)
@@
  • (C2)
@@
  • (C3)
    subject to (C1)-(C2), each vertex in $V$ lies in as many as possible $A_{i}\cup B_{i}$ for $i\in[2,m]$.
Denote $(A_{1},B_{1},L_{1}):=(\tilde{A}_{1},\tilde{B}_{1},\tilde{L}_{1})$. Expand $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$. For simplicity, we assume $Y\in\{A,B\}$ and $Z\in\{A,B\}\setminus\{Y\}$ in the following. Let $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ be a partition of $[m]$ with
[…]
Hence for $i\in[2,m]$, we have $|Y_{1}\Delta Y_{i}|\leq|\tilde{Y_{1}}\Delta\tilde{Y_{i}}|+|\tilde{L}_{i}|\leq\delta n+2{\epsilon}n<2\delta n$. Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise. Clearly, $|\hat{\mathcal{C}}|\geq(1-4\sqrt{\delta}-\eta)n$. Define
$X:=$$\{x\in V:x\not\in A_{i}\cup B_{i}\leavevmode\nobreak\ \textrm{forat least}\leavevmode\nobreak\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\},$
$X_{Y}:=$$\{x\in Y\setminus X:x\not\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 10\sqrt{\delta}|\hat{\mathcal{C}}|\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in\hat{\mathcal{C}}\}.$
Recall that $|L_{i}|\leq 2{\epsilon}n$ and $|Y_{1}\Delta Y_{i}|<2\delta n$ for each $i\in\hat{\mathcal{C}}$. Hence $6\sqrt{\delta}|\hat{\mathcal{C}}||X|\leq 2{\epsilon}n|\hat{\mathcal{C}}|$ and
$10\sqrt{\delta}|\hat{\mathcal{C}}||X_{Y}|\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq\sum_{i\in\hat{\mathcal{C}}}(|Y\setminus Y_{1}|+|Y_{1}\Delta Y_{i}|)\leq(2\delta+2{\epsilon})n|\hat{\mathcal{C}}|.$
$X_{Y}^{k}=\left\{\begin{array}[]{ll}\{x\in X_{Y}:x\in Y_{i}\leavevmode\nobreak\ \textrm{for at least}\leavevmode\nobreak\ 3\sqrt{\delta}n\leavevmode\nobreak\ \textrm{colors}\leavevmode\nobreak\ i\in{\mathcal{C}_{k}}\},&\textrm{if $|\mathcal{C}_{k}|\geq\eta n$,}\\[5.0pt] \emptyset,&\textrm{otherwise.}\end{array}\right.$
For a vertex $x\in X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})$, we know $x\in Y_{i}$ for at most $6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in A_{i}\cup B_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Thus, $x\in Y\setminus Y_{i}$ for at least $|\hat{\mathcal{C}}|-6\sqrt{\delta}n$ colors $i\in\hat{\mathcal{C}}$ and $x\in Z_{i}$ for at least $(1-6\sqrt{\delta})|\hat{\mathcal{C}}|-6\sqrt{\delta}n\geq(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Hence
$|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|(|\hat{\mathcal{C}}|-6\sqrt{\delta}n)\leq\sum_{i\in\hat{\mathcal{C}}}|Y\setminus Y_{i}|\leq(2\delta+2\epsilon)n|\hat{\mathcal{C}}|.$
It follows that $|X_{Y}\setminus(\bigcup_{k\in[2]}X_{Y}^{k})|<4\delta n$. Move vertices in $X_{A}\setminus(\bigcup_{k\in[2]}X_{A}^{k})$ to $B$ and vertices in $X_{B}\setminus(\bigcup_{k\in[2]}X_{B}^{k})$ to $A$.
Let $X^{\prime}$ be a subset of $X$ consisting of vertices $x$ such that $d_{i}^{+}(x,Y_{i})\geq(1-3\sqrt{\delta})n$ (resp. $d_{i}^{-}(x,Y_{i})\geq(1-3\sqrt{\delta})n$) for at least $(1-3\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$ and $d_{i}^{-}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$ (resp. $d_{i}^{+}(x,Z_{i})\geq\frac{5}{2}\sqrt{\delta}n$) for at least $3\sqrt{\delta}|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Delete vertices in $X^{\prime}$ from $A\cup B$. Without loss of generality, assume that $|B|-|A|=r$ with $0\leq r\leq 8{\delta}n+\sqrt{\epsilon}n$. Define $V_{\rm bad}:=\bigcup_{k\in[2]}(X^{k}_{A}\cup X^{k}_{B})\cup(X\setminus X^{\prime})$. Obviously, $|V_{\rm bad}|<(\sqrt{\epsilon}+\frac{1}{2}\sqrt{\delta})n$. Moreover, each vertex in $Y\setminus V_{\rm bad}$ lies in $Y_{i}$ for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$.
@@
The subsequent result can be used to connect two disjoint short rainbow directed paths into a single short rainbow directed path.
Claim 5.3 (Connecting tool).
@@
Assume $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s},v_{1}\notin V_{\rm bad}\cup X^{\prime}$ and $s+t\leq 5\eta n$.
@@
Note that $u_{s},v_{1}\not\in V_{\rm bad}\cup X^{\prime}$. Then $u_{s}$ (resp. $v_{1}$) lies in $Y_{i}$ (resp. $Z_{i}$) for at least $(1-13\sqrt{\delta})|\hat{\mathcal{C}}|$ colors $i\in\hat{\mathcal{C}}$. Since $|\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)|\geq 14\sqrt{\delta}n$, we obtain that $u_{s}$ (resp. $v_{1}$) lies in $Y_{i}$ (resp. $Z_{i}$) for at least $\sqrt{\delta}n$ colors $i\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$. Hence there are two distinct colors $c_{1},c_{2}\in\mathcal{C}_{2}\setminus{\rm col}(P\cup Q)$ such that $u_{s}\in Y_{c_{1}}$ and $v_{1}\in Z_{c_{2}}$. It is routine to check that
$|N_{{c_{1}}}^{+}(u_{s})\cap(Z\setminus(V_{\rm bad}\cup X^{\prime}))|\geq$ $|N_{{c_{1}}}^{+}(u_{s})\cap Z_{c_{1}}|-|Z_{1}\Delta Z_{c_{1}}|-|X\cup X_{A}\cup X_{B}|$

### v2

@@ 5 Extremal case
Assume that $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll 1$ and $m=(1-4\sqrt{\delta})n$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$. If $D_{i}$ is $(\epsilon,{\rm EC3})$-extremal for each $i\in[m]$, then $\mathcal{D}$ contains a transversal directed Hamilton cycle.
In the following two sections, we prove the above two theorems respectively.
6 Proof of Theorem 5.2
In this section we prove Theorem 5.2. Following the outline presented in Section 1.4, we first identify several types of “bad” vertices and cover them by short rainbow directed paths whose endpoints possess additional properties. Next we handle the “bad” colors by choosing matchings. We then connect all these short rainbow directed paths via a connecting lemma. Finally, by the transversal blow-up lemma, we construct a long transversal directed path, which can be closed to form the desired transversal directed cycle.
For the sake of clarity in presenting the following lemmas, we summarize some properties of the digraph collection $\mathcal{D}$.
  1. $(\dagger)$
    Let $0<\frac{1}{n}\ll\epsilon\leq\delta^{2}\ll\eta\ll 1$ and $m=(1-4\sqrt{\delta})n$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$ and $\delta(\mathcal{D})\geq\frac{n}{2}$. We assume the following conditions hold.
    1. (a)
      For every $i\in[m]$, $D_{i}$ is either $(\epsilon,{\rm EC1})$-extremal or $(\epsilon,{\rm EC2})$-extremal and it admits a partition $({A_{i}},{B_{i}},{L_{i}})$ satisfying (C1)-(C4). We extend $A_{1}\cup B_{1}$ to an equitable partition $A\cup B$ of $V$.
    2. (b)
      Define
      $\mathcal{C}_{1}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC1})\textrm{-extremal}\},$
      $\mathcal{C}_{2}=\{i\in[m]:D_{i}\penalty 10000\ \textrm{is}\penalty 10000\ (\epsilon,{\rm EC2})\textrm{-extremal}\},\ \ \text{and}\ \ \mathcal{C}_{\rm bad}=\mathcal{C}\setminus(\mathcal{C}_{1}\cup\mathcal{C}_{2}).$
      Let $\hat{\mathcal{C}}:=\bigcup_{k\in[2]}\psi(\mathcal{C}_{k}),$ where $\psi(\mathcal{C}_{k})=\mathcal{C}_{k}$ if $|\mathcal{C}_{k}|\geq\eta n$ and $\psi(\mathcal{C}_{k})=\emptyset$ otherwise.
    3. (c)
@@ 1. $(\dagger)$
    4. (d)
      Define
      $X:=\{x\in V:x\not\in A_{i}\cup B_{i}\penalty 10000\ \textrm{for at least}\penalty 10000\ 6\sqrt{\delta}|\hat{\mathcal{C}}|\penalty 10000\ \textrm{colors}\penalty 10000\ i\in\hat{\mathcal{C}}\}$.
@@
The next lemma shows that any vertex contained in $L_{i}$ for many colors $i\in\mathcal{C}$ can be covered by a rainbow directed path $P_{3}$ with endpoints not in $V_{\rm bad}$. To enhance flexibility, we select a rainbow star $S_{5}$ (instead of a simple directed $P_{3}$), which allows each endpoint of the resulting path to be replaced by an alternative available vertex when needed.
Lemma 6.1.
Suppose that $(\dagger)$6 holds and $\{Y,Z\}=\{A,B\}$. Let $P$ be a rainbow directed path in $\mathcal{D}$ with length at most $\frac{7}{3}\sqrt{\delta}n$. Assume $X\setminus(X^{\prime}\cup V(P))=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. If $x_{i}\in X\cap Y$, then there exist colors $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{j}\setminus({\rm col}(P)\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and vertices $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},x_{i}^{4}\in W\setminus(V(P)\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$, where
  • $j=2$ and $W=Z$ if $|\mathcal{C}_{1}|<\eta n$,
  • $j=1$ and $W=Y$ if $|\mathcal{C}_{2}|<\eta n$,
  • either $j=1$ and $W=Y$, or $j=2$ and $W=Z$ otherwise.
Proof.
We only prove the case that $|\mathcal{C}_{1}|<\eta n$, and the other cases can be proved by similar arguments. Suppose there exists an $i_{0}\in[s]$ such that Lemma 6.1 holds for all $i\in[i_{0}-1]$ but does not hold for $i_{0}$. Assume for convenience that $x_{i_{0}}\in X\cap A$. Denote $\mathcal{C}_{2}^{\prime}:=\mathcal{C}_{2}\setminus({\rm col}(P)\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i_{0}-1]\})$. Let
$\mathcal{C}_{2}^{1}:=\big\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{+}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P)|\big\}\ \text{and}$
$\mathcal{C}_{2}^{2}:=\big\{j\in\mathcal{C}_{2}^{\prime}:d_{j}^{-}(x_{i_{0}},B)\geq 4(i_{0}-1)+4+|V(P)|\big\}.$
$d_{j}^{+}(x_{i_{0}},A_{j})$ $\geq d_{j}^{+}(x_{i_{0}},A)-|A\setminus A_{1}|-|A_{1}\triangle A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\delta^{0}(D_{j})-d_{j}^{+}(x_{i_{0}},B)-|A\setminus A_{1}|-|A_{1}\triangle A_{j}|-|X_{A}\setminus X_{A}^{2}|-|X^{\prime}|$
$\geq\frac{n}{2}-\left(4(i_{0}-1)+4+|V(P)|\right)-2{\epsilon}n-2\delta n-4\delta n-\sqrt{\epsilon}n$
$\geq\left(\frac{1}{2}-3\sqrt{\delta}\right)n.$
Furthermore, $|\mathcal{C}_{2}^{\prime}\setminus\mathcal{C}_{2}^{1}|\geq(1-3\sqrt{\delta})|\mathcal{C}_{2}|$.
Notice that $x_{i_{0}}\notin X^{\prime}$. Then, for at least $(1-3\sqrt{\delta})|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, $d_{j}^{-}(x_{i_{0}},B_{j})<\frac{5}{2}\sqrt{\delta}n$, and consequently, $d_{j}^{-}(x_{i_{0}},A_{j})\geq(\frac{1}{2}-3\sqrt{\delta})n$ for the same set of colors. By (C4), we know that $x_{i_{0}}\in A_{j}\cup B_{j}$ for all but at most $6\sqrt{\delta}|\mathcal{C}_{2}|$ colors $j\in\mathcal{C}_{2}$, which leads to $x_{i_{0}}\not\in X$, a contradiction. ∎
Next, we give an application of the transversal blow-up lemma (see [10]) for embedding transversal undirected Hamilton paths inside very dense bipartite graph collections, which can be proved by minor modifications to the proof of [9, Lemma 6.1], and we omit the proof here. We say that $P$ is an undirected subgraph of a digraph collection $\mathcal{D}=\{{D_{1}},\ldots,D_{n}\}$ if, for every edge $uv\in E(P)\cap E(D_{i})$, the reverse arc $vu\in E(D_{i})$.
Lemma 6.2.
Suppose $(\dagger)$6 holds and $\{W,Z\}=\{A,B\}$. Let $W^{*}\subseteq W\setminus V_{\rm bad}$ and $Z^{*}\subseteq Z\setminus V_{\rm bad}$, where $|W^{*}|,|Z^{*}|\geq\eta n$, $W^{*}\cap Z^{*}=\emptyset$ and $|W^{*}|-|Z^{*}|=t\in\{0,1\}$. Let $T^{*}=Z^{*}$ if $t=0$ and $T^{*}=W^{*}$ if $t=1$. Let $\mathcal{C^{*}}\subseteq\mathcal{C}$ satisfy $|\mathcal{C^{*}}|=|W^{*}|+|Z^{*}|-1$, where $\mathcal{C^{*}}\subseteq\mathcal{C}_{1}$ if $W=Z$ and $\mathcal{C^{*}}\subseteq\mathcal{C}_{2}$ if $W\neq Z$. Let $W^{-}\subseteq W^{*}$ and $T^{+}\subseteq T^{*}$ with $|W^{-}|,|T^{+}|\geq\frac{\eta n}{8}$. Then there is a transversal undirected Hamilton path in $\{D_{i}^{\pm}[W^{*},Z^{*}]:i\in\mathcal{C}^{*}\}$ starting at $W^{-}$ and ending at $T^{+}$.
@@
The subsequent result can be used to connect two disjoint short rainbow paths into a single short rainbow path.
Lemma 6.3 (Connecting tool).
@@
Suppose that $(\dagger)$6 holds and $\{Y,Z\}=\{A,B\}$. Assume $P=u_{1}u_{2}\ldots u_{s}$ and $Q=v_{1}v_{2}\ldots v_{t}$ are two disjoint rainbow directed paths inside $\mathcal{D}$ with $u_{s},v_{1}\notin V_{\rm bad}\cup X^{\prime}$ and $s+t\leq 5\eta n$.
@@
[…]
$|N_{{c_{1}}}^{+}(u_{s})\cap(Z\setminus(V_{\rm bad}\cup X^{\prime}))|\geq$ $|N_{{c_{1}}}^{+}(u_{s})\cap Z_{c_{1}}|-|Z_{1}\triangle Z_{c_{1}}|-|X\cup X_{A}\cup X_{B}|$